OpenStudy (anonymous):

after doing a trig substitution and putting the original values back in what do i put in whenI have just theta

OpenStudy (jdoe0001):

what's the original issue?

OpenStudy (anonymous):

the original question was the indefinite integral of sqrt(2x-x^2)

OpenStudy (jdoe0001):

hmm... usually if you have \(\theta\) on one side, the other side will be the angle value for it

OpenStudy (anonymous):

yes I know but after substituting x=sin(theta) I end up with 1/2(2theta+sin(theta)(sqrt(-cos^2(theta))+theta(sqrt(-cos^2(theta))sec(theta)+c i can fin all the cos(theta), sin(theta) and sec(theta) by doing the triangle method its the thetas that are alone or multiplied by a constant that im nt sure how to handle

OpenStudy (jdoe0001):

hmm... dunno that one :(

OpenStudy (anonymous):

well thanks anyways :)

myininaya (myininaya):

if you made the sub tan(theta)=o/a then you say theta=arctan(o/a) or if you did the sub sin(theta)=o/h then you say theta=arcsin(o/h) if you did the sub cos(theta)=a/h then you say theta=arccos(a/h)

myininaya (myininaya):

for your case you used x=sin(theta) So replace theta with arcsin(x)